draw the graph of the linear equation 2x - 3y +7 = 0 and above hence find the coordinate of the point where the line intersect x axis

Answer by Alan3354(69443)

draw the graph of the linear equation 2x - 3y +7 = 0 and above hence find the coordinate of the point where the line intersect x axis
-----------------
Find the x-intercept first.
At the x-int, y = 0
2x - 3y +7 = 0
2x = -7
x = -7/2
---
Find the y-int where x = 0
2x - 3y +7 = 0
-3y = -7
y = 7/3
---
Plot the 2 points, (-7/2,0) and (0,7/3)
Draw a line thru them.